Two forces, F1 = 8.0 N, 60o Northwest and F2 = 12.0 N, Southwest, are applied to a particle at the origin. What third force is needed so that the particle will not be displaced from the origin?


1
Expert's answer
2021-02-15T00:55:40-0500

"-F_{3x}=F_1cos 60°+F_2cos45°,\\ F_{3x}=-12.49~\text{N},"

"F_{3y}=F_2sin45°-F_1sin60°,\\ F_{3y}=1.56~\text{N}."


"F=\sqrt{F_{3x}^2+F_{3y}^2}=12.59~\text{N},"

"\alpha=arctan|\frac{F_{3y}}{F_{3x}}|=83°~" East of North.